# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ B.4 Computational Complexity
We finally show that a tractable $\textsc{Cons}_{{\mathcal{F}}}$ oracle implies a computationally tractable ${\sf CoT}$ learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ via a chain-of-thought generated by ${\mathcal{F}}$. We first start with the proof of Theorem 3.6 which reduces ${\sf CoT}$-learnability to a consistency problem on the base class ${\mathcal{F}}$.
Proof of Theorem 3.6.
@@
It is easy to observe that if we find a consistent $\hat{f}\in{\mathcal{F}}$ on $\tilde{S}$ (in the sense of $\textsc{Cons}_{\mathcal{F}}$), then we also have that this predictor is consistent with the entire chain-of-thought, i.e. $\hat{f}^{{\sf CoT}\textnormal{-}{T}}({\bm{x}}_{i})={\bm{z}}_{i}$ for all $i\in[m]$. If the original inputs ${\bm{x}}_{i}$’s are of length at most $n$, then the prefixes $p_{(i,t)}$ are of length at most $n+T$. Moreover, we have $|\tilde{S}|=\tilde{m}\leq m\cdot T$ and thus, one such call of $\textsc{Cons}_{\mathcal{F}}$ on $\tilde{S}$ suffices. The initial input processing of creating $\tilde{S}$ from $S$, and also finally returning $\hat{f}^{{\sf e2e}\textnormal{-}{T}}$ can be done in additional time $O(\tilde{m})$, concluding the implementation of $\textsc{Cons}_{{\sf CoT}}$. ∎
Proof of Corollary 3.7.
This is a direct corollary of Theorems 3.6 and 3.4. If $\mathrm{VCdim}({\mathcal{F}}_{d})={\mathcal{P}oly}(d)$, then implementing $\textsc{Cons}_{{\sf CoT}}$ on a CoT training set $S_{{\sf CoT}}$ of size
$|S_{{\sf CoT}}|=O(\varepsilon^{\scriptscriptstyle{-1}}(\mathrm{VCdim}({\mathcal{F}}_{d})\log T\log\varepsilon^{\scriptscriptstyle{-1}}+\log\delta^{\scriptscriptstyle{-1}}))={\mathcal{P}oly}(d,T,\varepsilon^{\scriptscriptstyle{-1}},\delta^{\scriptscriptstyle{-1}})$
suffices for ${\sf CoT}$-learnability of ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ by Theorem 3.4. Moreover, by Theorem 3.6, we can implement this by calling $\textsc{Cons}_{{\mathcal{F}}_{d}}$ on a sample set $\tilde{S}$ of size at most $|\tilde{S}|=O(T\cdot|S_{{\sf CoT}}|)={\mathcal{P}oly}(d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. Finally, for input distributions that are supported on sequences of length at most $n$, the description length of $S_{{\sf CoT}}$ is ${\mathcal{P}oly}(n+T,d,|S_{{\sf CoT}}|)$. Here we are subsuming the $\log|\Sigma|$ dependence to be already captured in the size parameter as it is inherent to the base class ${\mathcal{F}}_{d}$. Thus even the description length of the set $\tilde{S}$ is ${\mathcal{P}oly}(n+T,d,|\tilde{S}|)={\mathcal{P}oly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. As such, as long as $\textsc{Cons}_{{\mathcal{F}}_{d}}$ is implementable in time polynomial in its input and the size parameter $d$, we have that ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf CoT}$-learnable in time ${\mathcal{P}oly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. ∎
@@ Appendix C Proofs from Section 4
Proof of Corollary 4.2.
The bound simply follows from Theorems 3.1, 3.2 and 3.4 after substituting $\mathrm{VCdim}(\mathcal{F}_{\scriptscriptstyle d,{\rm lin}})\leq(d+1)$ and $\log|\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}|=O(d^{2})$ by Lemma 4.1. ∎
@@
Proof of Corollary 4.3.
This directly follows from the facts (i) $\textsc{Cons}_{\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}}(S)$ is implementable in time polynomial in $d$ and the length of the input $S$ (ii) noting that $\mathrm{VCdim}(\mathcal{F}_{\scriptscriptstyle d,{\rm lin}})\leq(d+1)$. This is a corollary of Corollary 3.7.
@@
$\langle{\bm{w}},{\bm{p}}_{(i,t)}\rangle\geq\,$ 0,if $0,\quad\text{if ${\mathbf{y}}_{(i,t)}=1$;}$
$\langle{\bm{w}},{\bm{p}}_{(i,t)}\rangle<\,$ 0,if $0,\quad\text{if ${\mathbf{y}}_{(i,t)}=0$.}$ (LP-Feasibility)
Clearly there are $|S_{{\sf CoT}}|\cdot T$ many constraints, and we are implementing $\textsc{Cons}_{{\sf CoT}}$, and therefore choosing $|S_{{\sf CoT}}|={\mathcal{P}oly}(n,d,T,1/\varepsilon,\log(1/\delta))$ suffices by Corollary 4.2. The runtime complexity of solving (LP-Feasibility) with $d$ variables and $|S_{{\sf CoT}}|\times T$ constraints is bounded by ${\mathcal{P}oly}(n,d,T,1/\varepsilon,\log(1/\delta))$; the theorem follows. ∎
@@ C.1 End-to-End Learning is Hard for Iterated Linear Thresholds
For $n,L,s\in{\mathbb{N}}_{+}$, consider the hypothesis class threshold circuits over $n$ input variables, size $s$, and depth at most $L$. This class is referred in Assumption 1. We first prove the final hardness results in light of Lemma 4.5.
Proof of Theorem 4.4.
Consider the class threshold circuits of depth $L$ and size $p(n)$ from Hardness Assumption 1. By our expressivity result Lemma 4.5, every such circuit $C$ can be expressed as $f_{{\bm{w}}}^{{\sf e2e}\textnormal{-}{T}}$ with some ${\bm{w}}\in\mathbb{R}^{d}$ with $d\leq 2(p(n)+2)^{L}(n+1)$ and $T\leq(p(n)+2)^{L}(n+1)$, up to some fixed input transformation that runs in time $O(n^{\prime})=O\left(2(p(n)+2)^{L}(n+1)\right)$. All $d,n^{\prime},T$ are bounded by fixed polynomials in $n$. Therefore, if ${\mathcal{F}}_{\scriptscriptstyle{d,{\rm lin}}}$ is learnable in time ${\mathcal{P}oly}(n,T,d)$ up to even constants $\varepsilon,\delta>0$, we can also learn the class from Assumption 1 in time ${\mathcal{P}oly}(n)$ through this reduction, contradicting Assumption 1. ∎
@@
So roughly speaking, the size of each predictor $\tilde{{\bm{w}}}_{li}\in\mathbb{R}^{\tilde{p}_{l}}$ in layer $l$ that we embedded our ${\bm{w}}_{li}\in\mathbb{R}^{p_{l}}$ into, is one more than the length of the entire concatenation of vectors ${\bm{v}}_{l-1},\dots,{\bm{v}}_{0}$. And the vector ${\bm{v}}_{l}$ is a concatenation of $s$ such vectors, giving us $\tilde{p}_{l}=O(s^{l})$. See Figure 2.
Figure 2: The figure provides an illustrative summary of the construction so far and the sizes of the predictors, formalized in C.1.
@@
[…]
@@
where we used Claim C.1 and Eq. (30). Overall, our constructions so far satisfies the bounds on $T,n^{\prime},d$ according to Lemma 4.5. Therefore, it remains to show the final construction of $\tilde{{\bm{w}}}_{li}\in\mathbb{R}^{\tilde{p}_{l}}$ and ${\bm{v}}\in\mathbb{R}^{T}$, and that this construction achieves the desired end-to-end computation of the circuit. To this end, we first specify the set of time-steps at which we output a gate in the circuit. Let $t_{li}$ for $l\in[L]$ and $i\in[s]$ be the time when we are going to be writing $y_{li}$ (i.e. output of the $i^{\mathrm{th}}$ gate from the layer ${l}$). We define these time-steps recursively as
$t_{1i}=n\cdot i,\text{ for $i\in[s]$ }\quad\text{and}\quad t_{li}=t_{(l-1)s}+i\cdot\tilde{p}_{l}\;\;\text{ for $l>1,i\in[s]$ }\,.$ (35)
Let :=${\mathcal{I}}:=\{t_{li}:l\in[L],i\in[s]\}$ be the collection of these indices. It is straight-forward to verify that $t_{Ls}=\sum_{\ell=1}^{L}s\cdot\tilde{p}_{\ell}=\sum_{\ell=1}^{L}|{\bm{v}}_{\ell}|=T\,$ by Eq. (30). Therefore, the final output of the circuit $y_{Ls}$ will be computed at the $T^{\mathrm{th}}$ step (to be shown).
@@
This is exactly where the $1$ at the beginning of the feature map $\phi$ in Eq. (27) plays the role of “positional encoding” together with the vector ${\bm{v}}\in\mathbb{R}^{T}$. The main idea is to create a vector ${\bm{v}}\in\{-B,0\}^{T}$ where $-B$ is a “very large” negative value to be specified such that for every $t\in[T]\setminus{\mathcal{I}}$, the $1$ in $\phi({\bm{x}})$ aligns with $-B$ in ${\bm{v}}$, which on thresholding gives us the output 0. Formally,
${\bm{v}}[-t]:=\begin{cases}-B\,&,\text{if }t\in[T]\setminus{\mathcal{I}}\,;\\ \,\,0\,&,\text{if }t\in{\mathcal{I}}\,.\\ \end{cases}$ (37)
@@
[…]
@@
For every $l\in[L],i\in[s]$, we indeed have $|\tilde{{\bm{w}}}_{li}|=\tilde{p}_{l}$, and $f_{\bm{w}}^{{\sf e2e}\textnormal{-}{t_{li}}}(\phi({\bm{x}}))=y_{li}$ for $t_{li}\in{\mathcal{I}}$
@@
Proof of C.2.
We first note that during the output of the first $T$ predictions, our linear predictor has only shifted by at most $T-1$ positions to the right. At the beginning, the right-most coordinate ${\bm{v}}[-1]$ is aligned with $1$ to the left-end of $\phi({\bm{x}})$ according to the construction Eq. (33). As there are $T-1$ additional 00’s padded before ${\bm{x}}$, the weight ${\bm{v}}[-1]$ never gets aligned with any coordinates in ${\bm{x}}$ (or the new ones to be added to its right) in the first $T$ predictions.
In summary, letting ${\bm{z}}^{t}=f_{{\bm{w}}}^{{\sf CoT}\textnormal{-}{t}}(\phi({\bm{x}}))$, for any $11\leq t\leq T-$ 1, we can say
@@
where we used the fact that only weights from $[{\bm{v}}_{1},\dots,{\bm{v}}_{L}]$ align with $[{\bm{x}},{\bm{z}}^{t}[-t:]]$ during the first $T$ iterative steps. And these vectors are only constructed from $\{{\bm{w}}_{li}:l\in[L],i\in[s]\}$ by padding zeros, and ${\bm{z}}^{t}$ has binary coordinates. Hence, the inner-product is bounded by the total $\ell_{1}$ norm of the circuit.
Finally, during the $t^{\mathrm{th}}$ iterative prediction for any $t\in[T]\setminus{\mathcal{I}}$, we have that the coordinate of ${\bm{v}}$ that aligns with $1$ to the left in $\phi({\bm{x}})$ is ${\bm{v}}[-t]$; this is because ${\bm{v}}[-1]$ is aligned at the start and the predictor moved $t-1$ steps to the right. Therefore, using Eq. (39), for any $t\in[T]\setminus{\mathcal{I}}$,
@@
We first start by inductively (on $l$) verifying that $\tilde{{\bm{w}}}_{li}=\tilde{p}_{l}$ for $l\in[L],i\in[s]$. Clearly, by our description before (38), we don’t insert any zeros in ${\bm{w}}_{1i}$ for $1\leq i\leq s$, and therefore $|\tilde{{\bm{w}}}_{1i}|=p_{1}=n=\tilde{p}_{1}$. Now for any $l>1$, on the top of weights from the first $(l-2)$ layers, there are also $s$ number of weights associated with $(l-1)^{\mathrm{th}}$ layer in ${\bm{w}}_{li}$. Before all of these weights we have padded $\tilde{p}_{l-1}-1$ zeros. Therefore, using this and the induction hypothesis
@@
The base case is $t_{11}=n$. That means $n-1$ steps have already elapsed. By Claim C.2, we only output 00 in them. The linear predictor also moved $n-1$ positions to the right, and ${\bm{v}}_{0}=0^{n-1}$ aligns with the last $n-1$ tokens of CoT. The tokens before that are ${\bm{x}}$ which align with ${\bm{v}}_{1}[-n:]=\tilde{{\bm{w}}}_{11}={\bm{w}}_{11}$. Therefore, the output
@@
    Again $[{\bm{v}}_{l-1},\dots,{\bm{v}}_{0}]$ align with the last $\tilde{p}_{\ell}-1$ zeros and does not contribute. The $(i-1)\tilde{p}_{l}$ many weights before that in ${\bm{w}}$ are ${\bm{v}}_{l}[-(i-1)\tilde{p}_{l}:]=[\tilde{{\bm{w}}}_{l(i-1)},\dots,\tilde{{\bm{w}}}_{l1}]$ by Eq (29). Note that the only variables with potentially non-zero values are at the indices which are multiples of $\tilde{p}_{l}$. These correspond to the last coordinates of each $\tilde{{\bm{w}}}_{l(i-1)},\dots,\tilde{{\bm{w}}}_{l1}$ respectively; the value of all these coordinates is zero by our assumption that the last weight coming from the previous layer is always 0. The CoT before this (from ${\bm{x}}$ and new $t_{(l-1)s}$ CoT tokens) is exactly of the form we described in Case 1. Moreover, this now aligns with the $\tilde{p}_{l}$ many coordinates prior to $\tilde{{\bm{w}}}_{l(i-1)}$, which by Eq. (29) is exactly $\tilde{{\bm{w}}}_{li}$. Again noticing the form $\tilde{{\bm{w}}}_{li}$, we conclude

### v2

@@ B.4 Computational Complexity
We finally show that a tractable $\textsc{Cons}_{{\mathcal{F}}}$ oracle implies a computationally tractable $\sf CoT$} learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ via a chain-of-thought generated by $\mathcal{F}$}. We first start with the proof of Theorem˜3.6 which reduces $\sf CoT$}-learnability to a consistency problem on the base class $\mathcal{F}$}.
Proof of Theorem˜3.6.
@@
It is easy to observe that if we find a consistent $\hat{f}\in{\mathcal{F}}$ on $\tilde{S}$ (in the sense of $\textsc{Cons}_{\mathcal{F}}$), then we also have that this predictor is consistent with the entire chain-of-thought, i.e. $\hat{f}^{{\sf CoT}\textnormal{-}{T}}({\bm{x}}_{i})={\bm{z}}_{i}$ for all $i\in[m]$. If the original inputs ${\bm{x}}_{i}$’s are of length at most $n$, then the prefixes $p_{(i,t)}$ are of length at most $n+T$. Moreover, we have $|\tilde{S}|=\tilde{m}\leq m\cdot T$ and thus, one such call of $\textsc{Cons}_{\mathcal{F}}$ on $\tilde{S}$ suffices. The initial input processing of creating $\tilde{S}$ from $S$, and also finally returning $\hat{f}^{{\sf e2e}\textnormal{-}{T}}$ can be done in additional time O(m~)O($\tilde${m}), concluding the implementation of $\textsc{Cons}_{{\sf CoT}}$. ∎
Proof of Corollary˜3.7.
This is a direct corollary of Theorems˜3.6 and 3.4. If $\mathrm{VCdim}({\mathcal{F}}_{d})={\mathcal{P}\hskip-1.0ptoly}(d)$, then implementing $\textsc{Cons}_{{\sf CoT}}$ on a CoT training set $S_{{\sf CoT}}$ of size
$|S_{{\sf CoT}}|=O(\varepsilon^{\scriptscriptstyle{-1}}(\mathrm{VCdim}({\mathcal{F}}_{d})\log T\log\varepsilon^{\scriptscriptstyle{-1}}+\log\delta^{\scriptscriptstyle{-1}}))={\mathcal{P}\hskip-1.0ptoly}(d,T,\varepsilon^{\scriptscriptstyle{-1}},\delta^{\scriptscriptstyle{-1}})$
suffices for $\sf CoT$}-learnability of ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ by Theorem˜3.4. Moreover, by Theorem˜3.6, we can implement this by calling $\textsc{Cons}_{{\mathcal{F}}_{d}}$ on a sample set $\tilde{S}$ of size at most $|\tilde{S}|=O(T\cdot|S_{{\sf CoT}}|)={\mathcal{P}\hskip-1.0ptoly}(d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. Finally, for input distributions that are supported on sequences of length at most $n$, the description length of $S_{{\sf CoT}}$ is ${\mathcal{P}\hskip-1.0ptoly}(n+T,d,|S_{{\sf CoT}}|)$. Here we are subsuming the $\log|\Sigma|$ dependence to be already captured in the size parameter as it is inherent to the base class ${\mathcal{F}}_{d}$. Thus even the description length of the set $\tilde{S}$ is ${\mathcal{P}\hskip-1.0ptoly}(n+T,d,|\tilde{S}|)={\mathcal{P}\hskip-1.0ptoly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. As such, as long as $\textsc{Cons}_{{\mathcal{F}}_{d}}$ is implementable in time polynomial in its input and the size parameter $d$, we have that ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf CoT$}-learnable in time ${\mathcal{P}\hskip-1.0ptoly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. ∎
@@ Appendix C Proofs from Section 4
Proof of Corollary˜4.2.
The bound simply follows from Theorems˜3.1, 3.2 and 3.4 after substituting $\mathrm{VCdim}(\mathcal{F}_{\scriptscriptstyle d,{\rm lin}})\leq(d+1)$ and $\log|\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}|=O(d^{2})$ by Lemma˜4.1. ∎
@@
Proof of Corollary˜4.3.
This directly follows from the facts (i) $\textsc{Cons}_{\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}}(S)$ is implementable in time polynomial in $d$ and the length of the input $S$ (ii) noting that $\mathrm{VCdim}(\mathcal{F}_{\scriptscriptstyle d,{\rm lin}})\leq(d+1)$. This is a corollary of Corollary˜3.7.
@@
$\langle{\bm{w}},{\bm{p}}_{(i,t)}\rangle\geq\,$ $0,\quad\text{if ${\mathbf{y}}_{(i,t)}=1$;}$
$\langle{\bm{w}},{\bm{p}}_{(i,t)}\rangle<\,$ $0,\quad\text{if ${\mathbf{y}}_{(i,t)}=0$.}$ (LP-Feasibility)
Clearly there are $|S_{{\sf CoT}}|\cdot T$ many constraints, and we are implementing $\textsc{Cons}_{{\sf CoT}}$, and therefore choosing $|S_{{\sf CoT}}|={\mathcal{P}\hskip-1.0ptoly}(n,d,T,1/\varepsilon,\log(1/\delta))$ suffices by Corollary˜4.2. The runtime complexity of solving (LP-Feasibility) with $d$ variables and $|S_{{\sf CoT}}|\times T$ constraints is bounded by ${\mathcal{P}\hskip-1.0ptoly}(n,d,T,1/\varepsilon,\log(1/\delta))$; the theorem follows. ∎
@@ C.1 End-to-End Learning is Hard for Iterated Linear Thresholds
For $n,L,s\in{\mathbb{N}}_{+}$, consider the hypothesis class threshold circuits over $n$ input variables, size $s$, and depth at most $L$. This class is referred in Assumption˜1. We first prove the final hardness results in light of Lemma˜4.5.
Proof of Theorem˜4.4.
Consider the class threshold circuits of depth $L$ and size $p(n)$ from Hardness Assumption˜1. By our expressivity result Lemma˜4.5, every such circuit $C$ can be expressed as $f_{{\bm{w}}}^{{\sf e2e}\textnormal{-}{T}}$ with some ${\bm{w}}\in\mathbb{R}^{d}$ with $d\leq 2(p(n)+2)^{L}(n+1)$ and $T\leq(p(n)+2)^{L}(n+1)$, up to some fixed input transformation that runs in time $O(n^{\prime})=O\left(2(p(n)+2)^{L}(n+1)\right)$. All $d,n^{\prime},T$ are bounded by fixed polynomials in $n$. Therefore, if ${\mathcal{F}}_{\scriptscriptstyle{d,{\rm lin}}}$ is learnable in time ${\mathcal{P}\hskip-1.0ptoly}(n,T,d)$ up to even constants $\varepsilon,\delta>0$, we can also learn the class from Assumption˜1 in time ${\mathcal{P}\hskip-1.0ptoly}(n)$ through this reduction, contradicting Assumption˜1. ∎
@@
So roughly speaking, the size of each predictor $\tilde{{\bm{w}}}_{li}\in\mathbb{R}^{\tilde{p}_{l}}$ in layer $l$ that we embedded our ${\bm{w}}_{li}\in\mathbb{R}^{p_{l}}$ into, is one more than the length of the entire concatenation of vectors ${\bm{v}}_{l-1},\dots,{\bm{v}}_{0}$. And the vector ${\bm{v}}_{l}$ is a concatenation of $s$ such vectors, giving us $\tilde{p}_{l}=O(s^{l})$. See Figure˜2.
Figure 2: The figure provides an illustrative summary of the construction so far and the sizes of the predictors, formalized in ˜C.1.
@@
${\bm{w}}$ $=[\text{------}\,{\bm{v}}\,\text{------},\underbrace{{\bm{v}}_{L},{\bm{v}}_{L-1},\dots,{\bm{v}}_{2},{\bm{v}}_{1}}_{T\text{ coordinates}}\,\,,0^{n-1}]$
@@
where we used Claim C.1 and Eq. (30). Overall, our constructions so far satisfies the bounds on $T,n^{\prime},d$ according to Lemma 4.5. Therefore, it remains to show the final construction of $\tilde{{\bm{w}}}_{li}\in\mathbb{R}^{\tilde{p}_{l}}$ and ${\bm{v}}\in\mathbb{R}^{T}$, and that this construction achieves the desired end-to-end computation of the circuit. To this end, we first specify the set of time-steps at which we output a gate in the circuit. Let $t_{li}$ for $l\in[L]$ and $i\in[s]$ be the time when we are going to be writing $y_{li}$ (i.e. output of the $i^{\mathrm{th}}$ gate from the layer l{l}). We define these time-steps recursively as
[…]
Let ${\mathcal{I}}:=\{t_{li}:l\in[L],i\in[s]\}$ be the collection of these indices. It is straight-forward to verify that $t_{Ls}=\sum_{\ell=1}^{L}s\cdot\tilde{p}_{\ell}=\sum_{\ell=1}^{L}|{\bm{v}}_{\ell}|=T\,$ by Eq. (30). Therefore, the final output of the circuit $y_{Ls}$ will be computed at the $T^{\mathrm{th}}$ step (to be shown).
@@
This is exactly where the $1$ at the beginning of the feature map $\phi$ in Eq. (27) plays the role of “positional encoding” together with the vector ${\bm{v}}\in\mathbb{R}^{T}$. The main idea is to create a vector ${\bm{v}}\in\{-B,0\}^{T}$ where B-B is a “very large” negative value to be specified such that for every $t\in[T]\setminus{\mathcal{I}}$, the $1$ in $\phi({\bm{x}})$ aligns with B-B in $\bm{v}$}, which on thresholding gives us the output 0. Formally,
[…]
@@
 $\text{ for every }t\in[T]\setminus{\mathcal{I}}\,,\text{ we have }\;f_{{\bm{w}}}^{{\sf e2e}\textnormal{-}{t}}(\phi({\bm{x}}))=0\,.$
@@
For every $l\in[L],i\in[s]$, we indeed have $|\tilde{{\bm{w}}}_{li}|=\tilde{p}_{l}$, and $f_{\bm{w}}^{{\sf e2e}\textnormal{-}{t_{li}}}(\phi({\bm{x}}))=y_{li}$ for $t_{li}\in{\mathcal{I}}$ .
@@
Proof of ˜C.2.
We first note that during the output of the first $T$ predictions, our linear predictor has only shifted by at most T1T-1 positions to the right. At the beginning, the right-most coordinate ${\bm{v}}[-1]$ is aligned with $1$ to the left-end of $\phi({\bm{x}})$ according to the construction Eq. (33). As there are T1T-1 additional 0’s padded before $\bm{x}$}, the weight ${\bm{v}}[-1]$ never gets aligned with any coordinates in $\bm{x}$} (or the new ones to be added to its right) in the first $T$ predictions.
In summary, letting ${\bm{z}}^{t}=f_{{\bm{w}}}^{{\sf CoT}\textnormal{-}{t}}(\phi({\bm{x}}))$, for any $1\leq t\leq T-1$, we can say
@@
where we used the fact that only weights from ${\bm{v}}_{1},\dots,{\bm{v}}_{L}]$ align with ${\bm{x}},{\bm{z}}^{t}[-t:]]$ during the first $T$ iterative steps. And these vectors are only constructed from $\{{\bm{w}}_{li}:l\in[L],i\in[s]\}$ by padding zeros, and ${\bm{z}}^{t}$ has binary coordinates. Hence, the inner-product is bounded by the total $\ell_{1}$ norm of the circuit.
Finally, during the $t^{\mathrm{th}}$ iterative prediction for any $t\in[T]\setminus{\mathcal{I}}$, we have that the coordinate of $\bm{v}$} that aligns with $1$ to the left in $\phi({\bm{x}})$ is ${\bm{v}}[-t]$; this is because ${\bm{v}}[-1]$ is aligned at the start and the predictor moved t1t-1 steps to the right. Therefore, using Eq. (39), for any $t\in[T]\setminus{\mathcal{I}}$,
@@
We first start by inductively (on $l$) verifying that $\tilde{{\bm{w}}}_{li}=\tilde{p}_{l}$ for $l\in[L],i\in[s]$. Clearly, by our description before (38), we don’t insert any zeros in ${\bm{w}}_{1i}$ for $1\leq i\leq s$, and therefore $|\tilde{{\bm{w}}}_{1i}|=p_{1}=n=\tilde{p}_{1}$. Now for any $l>1$, on the top of weights from the first (l2)(l-2) layers, there are also $s$ number of weights associated with $(l-1)^{\mathrm{th}}$ layer in ${\bm{w}}_{li}$. Before all of these weights we have padded $\tilde{p}_{l-1}-1$ zeros. Therefore, using this and the induction hypothesis
@@
The base case is $t_{11}=n$. That means n1n-1 steps have already elapsed. By Claim C.2, we only output 0 in them. The linear predictor also moved n1n-1 positions to the right, and ${\bm{v}}_{0}=0^{n-1}$ aligns with the last n1n-1 tokens of CoT. The tokens before that are $\bm{x}$} which align with ${\bm{v}}_{1}[-n:]=\tilde{{\bm{w}}}_{11}={\bm{w}}_{11}$. Therefore, the output
@@
    Again ${\bm{v}}_{l-1},\dots,{\bm{v}}_{0}]$ align with the last $\tilde{p}_{\ell}-1$ zeros and does not contribute. The $(i-1)\tilde{p}_{l}$ many weights before that in $\bm{w}$} are ${\bm{v}}_{l}[-(i-1)\tilde{p}_{l}:]=[\tilde{{\bm{w}}}_{l(i-1)},\dots,\tilde{{\bm{w}}}_{l1}]$ by Eq (29). Note that the only variables with potentially non-zero values are at the indices which are multiples of $\tilde{p}_{l}$. These correspond to the last coordinates of each $\tilde{{\bm{w}}}_{l(i-1)},\dots,\tilde{{\bm{w}}}_{l1}$ respectively; the value of all these coordinates is zero by our assumption that the last weight coming from the previous layer is always 0. The CoT before this (from $\bm{x}$} and new $t_{(l-1)s}$ CoT tokens) is exactly of the form we described in Case 1. Moreover, this now aligns with the $\tilde{p}_{l}$ many coordinates prior to $\tilde{{\bm{w}}}_{l(i-1)}$, which by Eq. (29) is exactly $\tilde{{\bm{w}}}_{li}$. Again noticing the form $\tilde{{\bm{w}}}_{li}$, we conclude